# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ 00 $1$ $1$
We can show this using a Gröbner basis computation: Program [12, Program 6.6] computes the ideal generated by all the relations in (6.8) assuming $s_{1}=t_{1}=0$ and $s_{3}=t_{3}=0$, and checks that this is the whole ring. As a result, the system has no solution, and we either have $s_{3}\neq 0$ or $t_{3}\neq 0$. ∎
@@
so that the coefficient of $X^{2}$ in $P_{0}^{\prime}(X)$ is -1 as needed. Since $(Q_{1},Q_{2})$ is a prompt for a space $L_{15,2}$ and we applied a homothety to the zeroes of $Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}$, then by setting $Q_{1}^{\sharp}=a^{\sharp}P_{0}^{\sharp}$ for a suitable nonzero constant $a^{\sharp}$ and $Q_{2}^{\sharp}=P_{\infty}^{\sharp}$ we have that $(Q_{1}^{\sharp},Q_{2}^{\sharp})$ is a prompt for a space $L_{15,2}$ that is equivalent to $\Omega$. In particular, the coefficients of $Q_{1}^{\sharp}$ and $Q_{2}^{\sharp}$ satisfy the system (6.8). ∎
@@
$Q_{2}:=X^{5}+t_{2}X^{3}-t_{3}X^{2}+t_{4}X-t_{5}$
which is a prompt for a space $L_{15,2}$, denoted by $\Omega$. For every $\alpha\in\{-1,\frac{1}{a+1},-\frac{1}{a+1},\frac{1}{a-1},-\frac{1}{a-1}\}$ there exists a pair $(Q_{1}^{\sharp},Q_{2}^{\sharp})$ which is a prompt for the same space $\Omega$ such that $Q_{2}^{\sharp}$ is monic and $Q_{1}^{\sharp}=\alpha Q_{1}$.
@@
By Definition 3.3, there exists a constant $c\in k^{\times}$ such that the space $\Omega$ is generated by $\omega_{1}$ and $\omega_{2}$ with $\omega_{1}:=\frac{dX}{c^{3}(Q_{1}^{2}Q_{2}-Q_{2}^{3})}$ and $\omega_{2}:=\frac{dX}{c^{3}(Q_{1}^{3}-Q_{1}Q_{2}^{2})}$. We consider the following situations:
@@
    Let $(\omega_{1},-\omega_{2})$ be another basis of $\Omega$ and pick a pair $(Q_{1}^{\sharp},Q_{2}^{\sharp})$ associated with it. Since $-\omega_{2}$ has the same poles as $\omega_{2}$, then $Q_{2}^{\sharp}=Q_{2}$. We see then that $Q_{1}^{\sharp}=-Q_{1}$.
@@
    Let $(\omega_{1},\omega_{1}+\omega_{2})$ be another basis of $\Omega$ and pick a pair $(Q_{1}^{\sharp},Q_{2}^{\sharp})$ associated with it. By looking at the poles of $\omega_{1}+\omega_{2}$ we see that $Q_{2}^{\sharp}=\frac{Q_{1}+Q_{2}}{a+1}$. The leading coefficient of $Q_{1}^{\sharp}$ is computed as the ratio of leading coefficient of $Q_{1}$ by the one of $Q_{1}+Q_{2}$. In other words, $Q_{1}^{\sharp}=\frac{Q_{1}}{a+1}$.
The coefficients $-\frac{1}{a+1},\frac{1}{a-1},$ and $-\frac{1}{a-1}$ can be obtained by composing the changes of basis of $(i)$ and $(ii)$. ∎
Remark 6.12.
The transformations of the leading coefficients described in Lemma 6.11 are the unique possible if we want to keep the simplifications made previously. In fact, we are not allowed to change the zeroes of $A$ as this would change the parameter $s_{3}$ in the general case.
@@
Using the simplifications above, we are now ready to prove the main theorem. More precisely, we use Lemma 6.10 to assume that $s_{1}=t_{1}=0$ and $s_{3}=1$. Then, the program [12, Program 6.7] tells us that any solution to the system (6.8) needs to satisfy the following condition on the parameter $a$:
@@ 3. Case 3:
In all these cases, the successive elements of a Gröbner basis for the system (6.8) can be computed, and show us all the possibilities for the coefficients of $Q_{1}$ and $Q_{2}$. Let $\Delta_{2}(Q_{1},Q_{2})=Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}$ be the Moore determinant of $(Q_{1},Q_{2})$. We remark that, if $(Q_{1},Q_{2})$ and $(S_{1},S_{2})$ are prompts for spaces $L_{15,2}$ and $\Delta_{2}(Q_{1},Q_{2})=\Delta_{2}(S_{1},S_{2})$, then they actually are prompts for the same space by virtue of Corollary 2.10. Let us denote by $\Delta^{[1]}$ the Moore determinant of the pair $(Q_{1},Q_{2})$ appearing in Example 6.5, by $\Delta^{[2]}$ the Moore determinant of the pair $(Q_{1},Q_{2})$ appearing in Example 6.6, and by $\Delta^{\bullet}$ with an appropriate superscript the Moore determinants appearing in the cases below (e.g. $\Delta^{1.A}$ is the Moore determinant of the polynomials given by the program in case $1.A$). We show that we can get a complete classification up to equivalence by studying the cases outlined above and comparing the respective Moore determinants.
In Case 1, we let $a$ be a square root of $-1$. For simplicity and consistency, we assume that $a=-\mu^{20}$ where $\mu$ is the generator of ${\mathbb{F}}_{81}^{\times}$ appearing in Example 6.6. Then, the program [12, Program 6.8] returns the following subcases:
Case $1.A$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 0 1 $a-1$ 0 $-a+1$ $a$ $-1$ $a+1$
Case $1.A^{\prime}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 0 1 $-a-1$ 0 $a+1$ $-a$ $-1$ $-a+1$
@@
Case $1.B$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $-1$ 1 $a$ $1+a$ $-a$ 0 $a-1$ $a-1$
Case $1.B^{\prime}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $-1$ 1 $-a$ $1-a$ $a$ 0 $-(a+1)$ $-(a+1)$
@@
Case $1.C$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $a-1$ 1 $-1$ $a-1$ 0 $-a$ $a-1$ 0
Case $1.C^{\prime}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $-(a+1)$ 1 $-1$ $-(a+1)$ 0 $a$ $-(a+1)$ 0
@@
In Case 2, let $a$ be a root of $X^{3}-X^{2}+1$. For simplicity and consistency, we assume that $a=\mu^{-1}$ where $\mu$ is the generator of ${\mathbb{F}}_{27}^{\times}$ appearing in Example 6.5. Then we have that $\Phi(a)=\frac{1}{\mu+1}=-a^{2}-a-1$ and $\Phi^{2}(a)=\frac{1}{\mu-1}=a^{2}-1$. Program [12, Program 6.9] returns the following subcases:
Case $2.A$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $-1$ 1 $a^{2}$ $-(a^{2}+a+1)$ $-(a+1)$ $a^{2}+a+1$ $a^{2}-a$ $-(a^{2}-a+1)$
Case $2.B$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $-1$ 1 $-(a^{2}+a)$ $a$ $-(a+1)$ $a^{2}$ $-(a^{2}-a)$ 00
Case $2.C$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $-1$ 1 $a+1$ $a^{2}-1$ $-(a+1)$ $a^{2}-1$ $1$ $a^{2}-a-1$
@@
Finally, in Case 3 let $a$ be a root of $X^{3}-X+1$. For simplicity and consistency, we assume that $a=\mu$, the generator of ${\mathbb{F}}_{27}^{\times}$ appearing in Example 6.5. Then we have that $\Phi(a)=a-1$ and $\Phi^{2}(a)=a+1$. Program [12, Program 6.10] returns the following subcases:
Case $3.A$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $a$ 1 1 $-(a^{2}-a+1)$ $-(a^{2}+a+1)$ $-(a+1)$ $a^{2}+a$ $-(a^{2}+a+1)$
Case $3.B$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $-(a^{2}+a)$ 1 $a^{2}-a$ $a$ $-a^{2}$ $a-1$ $a^{2}-1$ $a^{2}$
Case $3.C$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $-(a^{2}-a+1)$ 1 $a^{2}+1$ 00 $-(a^{2}+a)$ $a^{2}$ $-a$ $a^{2}+a-1$

### v2

@@ 6.2. Classification of spaces $L_{15,2}$
We can show this using a Gröbner basis computation: Program [GithubRepo, Program 6.6] computes the ideal generated by all the relations in (6.8) with $s_{1}=t_{1}=0$ and $s_{3}=t_{3}=0$, and verifies that the system (6.8) has no solution under these assumptions. As a result, we either have $s_{3}\neq 0$ or $t_{3}\neq 0$. ∎
@@
so that the coefficient of $X^{2}$ in $P_{0}^{\prime}(X)$ is -1 as needed. Since $(Q_{1},Q_{2})$ is a prompt for a space $L_{15,2}$ and we applied a homothety to the zeroes of $Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}$, then by setting $Q_{1}^{\sharp}=P_{\infty}^{\sharp}$ and $Q_{2}^{\sharp}=a^{\sharp}P_{0}^{\sharp}$ for a suitable nonzero constant $a^{\sharp}$ we have that $(Q_{1}^{\sharp},Q_{2}^{\sharp})$ is a prompt for a space $L_{15,2}$ that is equivalent to $\Omega$. In particular, the coefficients of $Q_{1}^{\sharp}$ and $Q_{2}^{\sharp}$ satisfy the system (6.8). ∎
@@
$Q_{2}:=a(X^{5}+s_{2}X^{3}-X^{2}+s_{4}X-s_{5})$
which is a prompt for a space $L_{15,2}$, denoted by $\Omega$. For every $\alpha\in\{-1,\frac{1}{a+1},-\frac{1}{a+1},\frac{1}{a-1},-\frac{1}{a-1}\}$ there exists a pair $(Q_{1}^{\sharp},Q_{2}^{\sharp})$ which is a prompt for the same space $\Omega$ such that $Q_{1}^{\sharp}$ is monic and $Q_{2}^{\sharp}=\alpha Q_{2}$.
@@
By Definition 3.5, there exists a constant $c\in k^{\times}$ such that the space $\Omega$ is generated by $\omega_{1}$ and $\omega_{2}$ with $\omega_{1}:=\frac{dX}{c^{3}(Q_{1}Q_{2}^{2}-Q_{1}^{3})}$ and $\omega_{2}:=\frac{dX}{c^{3}(Q_{1}^{2}Q_{2}-Q_{2}^{3})}$. We first consider the following situations:
@@
    If we consider the basis $(\omega_{1},-\omega_{2})$ of $\Omega$ and apply Proposition 4.1.i, we find that the pair $(Q_{1},-Q_{2})$ is a prompt for $\Omega$.
@@
    If we consider the basis $(\omega_{1},\omega_{1}+\omega_{2})$ of $\Omega$ and apply Proposition 4.1.i, we find that the pair $(Q_{1}-Q_{2},Q_{2})$ is a prompt for $\Omega$. Hence, the pair $(\frac{Q_{1}-Q_{2}}{1-a},\frac{Q_{2}}{1-a})$ is also a prompt for $\Omega$ and its first term is monic.
The coefficients $\frac{1}{a-1},\frac{1}{a+1},$ and $\frac{1}{a+1}$ can be obtained similarly by considering all the other possible bases of $\Omega$. ∎
@@
Using the simplifications above, we are now ready to prove the main theorem. More precisely, we use Lemma 6.10 to assume that $s_{1}=t_{1}=0$ and $s_{3}=1$. Then, the program [GithubRepo, Program 6.7] tells us that any solution to the system (6.8) needs to satisfy the following condition on the parameter $a$:
@@ 3. Case 3:
In all these cases, the successive elements of a Gröbner basis for the system (6.8) can be computed, and show us all the possibilities for the coefficients of $Q_{1}$ and $Q_{2}$. Let $\Delta_{2}(Q_{1},Q_{2})=Q_{1}Q_{2}^{3}-Q_{1}^{3}Q_{2}$ be the Moore determinant of $(Q_{1},Q_{2})$. We remark that, if $(Q_{1},Q_{2})$ and $(S_{1},S_{2})$ are prompts for spaces $L_{15,2}$ and $\Delta_{2}(Q_{1},Q_{2})=\Delta_{2}(S_{1},S_{2})$, then these pairs are prompts for the same space by virtue of Corollary 2.11. With this in mind, we denote by $\Delta^{[1]}$ the Moore determinant of the pair $(Q_{1},Q_{2})$ appearing in Example 6.5, by $\Delta^{[2]}$ the Moore determinant of the pair $(Q_{1},Q_{2})$ appearing in Example 6.6, and by $\Delta^{\bullet}$ with an appropriate superscript the Moore determinants appearing in the cases below (e.g. $\Delta^{1.A}$ is the Moore determinant of the polynomials given by the program in case $1.A$). We show that we can get a complete classification up to equivalence by studying the cases outlined above and comparing the respective Moore determinants.
In Case 1, we let $a$ be a square root of 1-1. For simplicity and consistency, we assume that $a=-\mu^{20}$ where $\mu$ is the generator of ${\mathbb{F}}_{81}^{\times}$ appearing in Example 6.6. Then, the program [GithubRepo, Program 6.8] returns the following subcases:
Case $1.A$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 0 1 a1a-1 0 a+1-a+1 $a$ 1-1 $a+1$
Case $1.A^{\prime}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 0 1 a1-a-1 0 $a+1$ a-a 1-1 a+1-a+1
@@
Case $1.B$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 1-1 1 $a$ $1+a$ a-a 0 a1a-1 a1a-1
Case $1.B^{\prime}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 1-1 1 a-a 1a1-a $a$ 0 (a+1)-(a+1) (a+1)-(a+1)
@@
Case $1.C$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ a1a-1 1 1-1 a1a-1 0 a-a a1a-1 0
Case $1.C^{\prime}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ (a+1)-(a+1) 1 1-1 (a+1)-(a+1) 0 $a$ (a+1)-(a+1) 0
@@
In Case 2, let $a$ be a root of $X^{3}-X^{2}+1$. For simplicity and consistency, we assume that $a=\mu^{-1}$ where $\mu$ is the generator of ${\mathbb{F}}_{27}^{\times}$ appearing in Example 6.5. Then we have that $\Phi(a)=\frac{1}{\mu+1}=-a^{2}-a-1$ and $\Phi^{2}(a)=\frac{1}{\mu-1}=a^{2}-1$. Program [GithubRepo, Program 6.9] returns the following subcases:
Case $2.A$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 1-1 1 $a^{2}$ $(a^{2}+a+1)$ (a+1)-(a+1) $a^{2}+a+1$ $a^{2}-a$ $(a^{2}-a+1)$
Case $2.B$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 1-1 1 $(a^{2}+a)$ $a$ (a+1)-(a+1) $a^{2}$ $(a^{2}-a)$ 0
Case $2.C$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ 1-1 1 $a+1$ $a^{2}-1$ (a+1)-(a+1) $a^{2}-1$ 11 $a^{2}-a-1$
@@
Finally, in Case 3 let $a$ be a root of $X^{3}-X+1$. For simplicity and consistency, we assume that $a=\mu$, the generator of ${\mathbb{F}}_{27}^{\times}$ appearing in Example 6.5. Then we have that $\Phi(a)=a-1$ and $\Phi^{2}(a)=a+1$. Program [GithubRepo, Program 6.10] returns the following subcases:
Case $3.A$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $a$ 1 1 $(a^{2}-a+1)$ $(a^{2}+a+1)$ (a+1)-(a+1) $a^{2}+a$ $(a^{2}+a+1)$
Case $3.B$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $(a^{2}+a)$ 1 $^{2}-a aa$ $a^{2}$ a1a-1 $a^{2}-1$ $a^{2}$
Case $3.C$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$ $(a^{2}-a+1)$ 1 $a^{2}+1$ 0 $(a^{2}+a)$ $^{2} a-a$ $a^{2}+a-1$